OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. Round to the nearest tenth if necessary x^2+5x-84=0 {-12, 7} {7, 12} {-7, 12} No Real Solution

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

\[\Delta = b ^{2} -4ac = 25 - 4 (1)(-84)= 361\] \[x = (- b \pm \sqrt{\Delta}) / 2a = (-5 \pm \sqrt{61})/2(1) = (-5 \pm 19)/2(1)\] X= 7 or -12

OpenStudy (anonymous):

Or you could simply do like this: \[x ^{2} + 5x - 84 = x ^{2} - 7x + 12x - 84 = x(x-7) + 12(x-7) = (x-7)(x+12)\] Thus, x = 7 or -12